OpenStudy (anonymous):

]A 10−pound package of cookies costs (y + 1) dollars while a 6−pound package costs (y − 1) dollars. If the price per pound of cookies is the same for both packages, the equation below can be used to solve for y. All of y plus 1 divided by 10 equals all of y minus 1 divided by 6 What is the value, in dollars, of y? (Do not put a $ in your answer.)

OpenStudy (anonymous):

@insanitycat

OpenStudy (tkhunny):

The definitions are like this "(y+1) dollars" See how the "dollars" is listed explicitly? This means 'y' will not have that unit. The unit is already supplied. \(\dfrac{(y+1)\;dollars}{$10} = \dfrac{(y-1)\; dollars}{$6}\) \(\dfrac{y+1}{10} = \dfrac{y-1}{6}\) -- Observe that \(\dfrac{dollars}{$} = 1\) and we can simplify the equation by removing the unnecessary factor.